# Benchmarking Web API Integration Code Generation

### v5

@@ III-A Results
[…]
Figure 2: Performance comparison between selected models for full completion (top) and argument completion (bottom).
Figure 2 shows key performance metrics (explained in Table II) for one representative [^11] With a size as close as possible to 15B for a fair comparison. of each model family plus the best-performing open-source model for full and argument completion, respectively. The corresponding numbers are provided in Tables III and IV. We used greedy decoding for all experiments.
@@ III-B Discussion
The models perform comparatively well in predicting the HTTP method, which is expected since there is only a small set of available methods, with GET and POST being by far the most common ones (cf. Table I). In contrast, URLs and arguments are challenging to predict due to their unrestricted nature. The argument precision may be inflated to some extent by the Authorization header argument, which is required by most endpoints and therefore easy to predict. The relatively high argument value conditional accuracy indicates that identifying the correct argument is harder for the models than assigning the correct value to it once it is found.
@@
RQ: How correct is LLM-generated API invocation code, and what types of errors are commonly present in such code?
@@
Answer: While LLMs are able to generate partially correct API invocations, the overall result is incorrect most of the time. The types of errors commonly present are manifold and range from selecting the wrong endpoint to leaving out required arguments or hallucinating entirely illegal arguments.
@@
TABLE III: Evaluation Results for Full Completion
CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Code Llama (13B) Code Llama (70B) Llama 3.1 (8B) GPT-4o
Correct implementations (t) 000.06 000.12 000.26 000.18 000.21 000.00 000.08 000.30 000.00 000.60
Correct URLs (e) 000.32 000.54 000.60 000.60 000.53 000.00 000.58 000.72 000.00 000.78
Illegal URLs (e) 000.39 000.18 000.22 000.20 000.21 000.00 000.24 000.15 000.00 000.09
Correct methods (e) 000.85 000.88 000.87 000.88 000.84 000.00 000.84 000.89 000.00 000.91
Illegal methods (e) 000.13 000.09 000.11 000.11 000.14 000.00 000.14 000.10 000.00 000.07
Mean argument precision (e) 000.52 000.65 000.68 000.65 000.67 000.00 000.72 000.75 000.00 000.89
Mean argument recall (e) 000.43 000.54 000.63 000.56 000.57 000.00 000.65 000.69 000.00 000.87
Mean argument value conditional accuracy (e) 000.53 000.68 000.79 000.72 000.76 000.00 000.57 000.84 000.00 000.95
@@
TABLE IV: Evaluation Results for Argument Completion
CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Code Llama (13B) Code Llama (70B) Llama 3.1 (8B) GPT-4o
Correct implementations (t) 000.12 000.27 000.25 000.26 000.35 000.37 000.12 000.40 000.05 000.77
Illegal implementations (e) 000.63 000.27 000.36 000.33 000.32 000.35 000.30 000.23 000.31 000.09
Mean argument precision (e) 000.58 000.81 000.71 000.74 000.77 000.74 000.79 000.83 000.65 000.94
Mean argument recall (e) 000.53 000.74 000.65 000.67 000.71 000.72 000.74 000.79 000.43 000.94
Illegal arguments (e) 000.25 000.13 000.14 000.15 000.13 000.20 000.13 000.09 000.16 000.05
Mean argument value conditional accuracy (e) 000.71 000.83 000.83 000.83 000.86 000.78 000.65 000.89 000.67 000.95
III-C Limitations and Threats to Validity
@@
1) Real-world API integration code is diverse and complex. To be able to evaluate API invocations automatically, we had to apply some restrictions and simplifications when creating the tasks in our dataset, as described in Section II-A. For instance, there is limited program context before the API invocation. Additionally, we evaluate only the correctness of the outgoing request, not how the incoming response is handled. These limitations do, however, not interfere with our investigation of LLM’s ability to recall information about correct API usage—if this information is memorized, they should be able to recall it in our evaluation as well as real-world settings.
@@
2) The quality of the dataset and evaluation depends on the quality of the API specifications. We encountered and manually fixed several errors in the real-world specifications. Moreover, the specifications sometimes contain usage constraints that are only explained in a parameter’s free-form textual description and thus not captured by our automated correctness analysis.
@@
IV Related Work
@@
Since APIs are such a universal concept, it is important to distinguish between different settings in which APIs are being used and understand their different characteristics when discussing related work. We divide APIs roughly into the following categories: 1. general web APIs(this work), 2. domain-specific web APIs, 3. SDK-wrapped web APIs, 4. local function APIs, and 5. tool APIs for AI agents.
@@
2) Su et al. [27, 28] and Hosseini et al. [29] investigated natural language interfaces to domain-specific web APIs following the Open Data Protocol [^12] https://docs.oasis-open.org/odata/odata/v4.0/odata-v4.0-part1-protocol.html (OData). Their work has a narrow focus on data-centric applications (mostly within the Microsoft ecosystem) and an SQL-like syntax. In contrast, our benchmark targets OpenAPI, a very general and flexible API standard, making the task more open-ended and challenging. Also, instead of ad-hoc API queries by a user, we strive for production-ready API integration code implemented in JavaScript.
3) Works like Dialog2API [30] and CloudAPIBench [31] focused on SDK-wrapped web API invocations (e.g., for AWS), treating web APIs like local libraries. Both approaches do not generalize due to SDK heterogeneity. In contrast, our benchmark relies on a unified RESTful interface for any web API. This eliminates SDK dependency, enabling direct assessment of models’ core API comprehension rather than SDK-specific knowledge. By standardizing interactions across diverse APIs, we remove evaluation biases and isolate model capabilities from implementation quirks of individual SDKs.
4) Local function APIs are the most common kind of API in typical codebases and studied in many works, e.g., [32, 33, 34]. Invoking a function requires specifying the function name and one argument list. In contrast, general web API invocations are subject to more complex syntax and semantics than such local function calls, as already mentioned in Section I (Scope).
5) There is a large body of work concerned with enabling LLM-based AI agents to use tools to interact with their environment [35]. These tools range from simple utilities like calculators [36] to web services that enable AI agents to perform tasks such as weather queries or restaurant booking [37, 38, 39]. At first sight, this body of work seems related to research on automatic generation of integration code, as tools typically expose a function-like interface. However, both input and output modalities differ between these settings, as well as their non-functional requirements, so benchmarks are not directly transferable to other kinds of API invocations.
@@
Since functional abstractions are essential in programming, virtually every code dataset includes (function) API invocations. Code datasets involving web APIs can be found in the fields of natural language interfaces to web APIs [27, 29], SDK-based API invocations [30, 31], and tool-using AI agents [40, 37, 38, 39, 41]. We are not aware of datasets specifically about generating API integration code to invoke REST APIs like we do.
@@
The correctness of AI-generated code is traditionally measured either using exact matches or syntax-based similarity, which both correlate weakly with actual correctness [23], or using functional testing [24], which is not directly transferable to web API invocations, as discussed in Section II-C. This is why we developed a novel approach described in Section II. Evaluation methods in related works on API invocations mostly rely on variations of the aforementioned standard approaches. Some works use humans [42] or LLMs [39] as judges, and others execute the API invocation and verify the state of the environment [30, 41]. We are not aware of an evaluation method able to provide results as fine-grained as ours.
@@
LLMs display a remarkable capacity to memorize information from their training data—which can be beneficial when it comes to generating factual responses, but also a problem when it comes to leaking private information from the training data. Studies have shown that LLMs memorize, among other things, method definitions and method calls [43], but also that they struggle to suggest the correct APIs [11]. These findings are in line with our observation that LLMs seem to know many endpoint URLs and argument names, but are often unable to assemble them to a specification-compliant API invocation that solves the task in question.
If LLMs are unable to recall memorized facts, they tend to hallucinate information that is not grounded in reality. This is a common error of LLMs, and studies have shown that it also exists in code generation, where, among other things, models may hallucinate function names and arguments [44, 6, 7]. Our study confirms that these findings also hold for web API endpoints and their arguments, where a high amount of hallucinations can be observed.
@@
Correct usage of APIs is challenging also for humans, as indicated by several studies on API usability issues [3, 4] and real-world API misuses [45, 46, 47]. Several (static analysis) methods have been proposed for detecting (and mitigating) API misuses, e.g., [45, 48, 47]. They are, however, primarily designed to aid humans, not LLMs. Using static analyses for evaluation purposes is generally conceivable, but they can only detect whether an API invocation is legal or illegal, not whether it also solves the given task.
V Conclusion and Future Work

### v6

@@ 3.1. Results
Figure 2. Performance comparison between selected models for full completion (left) and argument completion (right) with unconstrained decoding
Figure 4 shows key performance metrics (explained in Table 2) for one representative [^11] With a size as close as possible to 15B for a fair comparison. of each model family plus the best-performing open-source model for full and argument completion, respectively. The corresponding numbers are provided in Tables 3 and 4. We used greedy decoding for all experiments.
@@ 3.2. Discussion
Table 3. Evaluation results for full completion with unconstrained decoding
CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Llama 3.1 (8B) Code Llama (13B) Code Llama (70B) GPT-4o
Correct implementations (t) 000.06 000.12 000.26 000.17 000.21 000.00 000.00 000.08 000.30 000.60
Correct URLs (e) 000.32 000.54 000.60 000.60 000.53 000.00 000.00 000.58 000.72 000.78
Illegal URLs (e) 000.39 000.18 000.22 000.20 000.21 000.00 000.00 000.24 000.15 000.09
Correct methods (e) 000.85 000.88 000.87 000.88 000.84 000.00 000.00 000.84 000.89 000.91
Illegal methods (e) 000.13 000.09 000.11 000.11 000.14 000.00 000.00 000.14 000.10 000.07
Mean argument precision (e) 000.53 000.67 000.68 000.66 000.70 000.00 000.00 000.73 000.75 000.89
Mean argument recall (e) 000.43 000.54 000.63 000.56 000.57 000.00 000.00 000.65 000.69 000.87
Mean arg. val. conditional accuracy (e)
000.72 000.84 000.92 000.88 000.91 000.00 000.00 000.63 000.91 000.96
Table 4. Evaluation results for argument completion with unconstrained decoding
CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Llama 3.1 (8B) Code Llama (13B) Code Llama (70B) GPT-4o
Correct implementations (t) 000.12 000.27 000.25 000.26 000.35 000.37 000.05 000.12 000.40 000.77
Illegal implementations (e) 000.63 000.27 000.36 000.33 000.32 000.35 000.31 000.30 000.23 000.09
Mean argument precision (e) 000.59 000.81 000.72 000.74 000.80 000.74 000.73 000.79 000.83 000.94
Mean argument recall (e) 000.53 000.74 000.65 000.67 000.71 000.72 000.43 000.74 000.79 000.94
Illegal arguments (e) 000.25 000.13 000.14 000.15 000.13 000.20 000.16 000.13 000.09 000.05
Mean arg. val. conditional accuracy (e)
000.86 000.85 000.90 000.91 000.93 000.89 000.89 000.70 000.92 000.97
The models perform comparatively well in predicting the HTTP method, which is expected since there is only a small set of available methods, with GET and POST being by far the most common ones (cf. Table 1). In contrast, URLs and arguments are challenging to predict due to their unrestricted nature. The argument precision may be inflated to some extent by the Authorization header argument, which is required by most endpoints and therefore easy to predict. The relatively high argument value conditional accuracy indicates that identifying the correct argument is harder for the models than assigning the correct value to it once it is found.
@@
RQ1: How correct is LLM-generated web API invocation code, and what types of errors are prevalent in such code? Answer: While LLMs are able to generate partially correct API invocations, the overall result is incorrect most of the time. The types of errors commonly present are manifold and range from selecting the wrong endpoint to leaving out required arguments or hallucinating entirely illegal arguments.
4. Translation of API Specifications to Constraints
Refer to caption Flow diagram visualizing the constrained decoding framework described in the caption.
Figure 3. Constrained decoding framework for generating web API invocations. In regular autoregressive generation, a decoder takes the token sequence $s_{1},\dots,s_{i}$ and predicts the probabilities $p_{j}$ for each possible next token $t_{j}$, based on which $s_{i+1}$ is selected. Constrained decoding augments this loop with a constraint check that sets the probability of tokens that fail the check to zero. We derive our constraints from OpenAPI specifications, and they ensure that generated API invocations comply with the specification.
Constrained decoding (Hokamp and Liu, 2017; Deutsch et al., 2019) is a general technique that modifies an LLM’s next-token predictions based on predefined constraints to prevent the generation of tokens that would lead to undesirable output sequences. As this technique appears to be underexplored for web API invocation code while also showing promise in related settings (cf. Section 7), we want to study its efficacy for web API invocation and its impact on the issues we uncovered in Section 3.
We implemented a simple, model-agnostic constrained decoding framework for our investigations, which is depicted in Figure 3. The constraints are automatically derived from OpenAPI specifications and represented as regular expressions (regex). They are designed to ensure that generated API invocations are always compliant with the API’s specification. Note that our contribution here lies in this transformation of specifications into constraints, and not the implementation of the constrained decoding engine. While we developed a custom implementation to maintain full control over the system and to add some debugging features, migrating to an off-the-shelf engine (cf. Section 7) would be feasible thanks to the engine-agnostic nature of our constraints; such a migration could enhance performance and robustness.
A main challenge when translating OpenAPI specifications to regular expressions in a generalizable and scalable way is the sheer size and complexity of real-world APIs, which is also reflected in the respective specification files. Each API can have dozens of endpoints (in the APIs we worked with, up to 174), and each endpoint can have dozens of parameters, distributed on different locations (body, header, path, query). Parameters can be optional or required, and they can be passed in arbitrary order. Additionally, many parameters have complex data types (objects and arrays) with a strictly defined nesting structure. There are also interdependencies between different parts of an API invocation, e.g., only for certain HTTP methods, parameters can be placed in the request body. In summary, some elements of an API invocations must be followed strictly, while other elements are flexible.
Additional challenges arise at the syntax-level of the language the API invocation is implemented in. Our regexes should accept valid JavaScript while not limiting variations permitted by the language, such as formatting or the use of different quotation marks. This is important as different LLMs show different stylistic preferences—forcing them away from their preferred style may degrade performance. There are also a lot of edge cases that need to be handled carefully by the constraints, e.g., an escaped quotation mark must not terminate a string literal, and many other seemingly minor details—our experiments showed that models will find and exploit such loopholes.
We address the challenges by structuring constraints hierarchically from the HTTP method and the URL to the different parameters. Once the method is matched, the search space is reduced to URLs for which an endpoint with the respective method is defined. Next, once the URL is matched, the search space is reduced to the parameters of the identified endpoint. Finally, once the parameter name is matched, the search space is reduced to values with the right data type. Within the parameter lists of a given endpoint, we use capture groups [^12] https://docs.python.org/3/library/re.html to keep track of parameters that have been defined. Conditional patternsfootnotemark: are then used to check the existence of a specific group and thus guard against the double definition of a parameter. Moreover, conditionals are used in conjunction with negative lookahead assertionsfootnotemark: to make it impossible to close a parameter list before all required parameters have been defined.
While we could represent all our constraints as one monolithic regex, we opted for a more modular and flexible approach [^13] They could still be joined if needed for compatibility with other frameworks.. We organize constraints as a set of so-called generation rules. Each generation rule has a regex-encoded start condition, stop condition, and body. Once the start condition matches the code generated so far, we constrain the generation according to the body until the stop condition matches.
There is one specialized generation rule for each HTTP method, which is triggered when the corresponding Axios invocation is generated (e.g., axios.post(). As multiple generation rules may be active at the same time (a token is only admissible if no rule rejects it), we can leverage overlapping constraints. We do so via a special generation rule that becomes active before all other rules and guides the generation towards the start condition of the other rules, ensuring one of them is going to be activated (thus avoiding using methods not supported by the given API). Moreover, we let generation rule bodies overlap with their stop conditions to ensure they properly terminate. This approach can also be generalized to many other applications of constrained decoding to orchestrate multiple constraints in a flexible and efficient manner.
5. Evaluation of Constrained LLMs
We put our constraints into action with all open-source models [^14] As they are closed-source, we cannot apply constrained decoding to the GPT models. under test and reuse WAPIIBench’s dataset and evaluation pipeline from Section 2. Again, we focus the analysis in this section on a subset of models, but provide comprehensive results in the appendix.
5.1. Results
[…]
Figure 4. Performance comparison between selected models for full completion (left) and argument completion (right) with constrained decoding
Refer to caption Stacked bar chart visualizing the performance difference between unconstrained and constrained decoding.
Figure 5. Relative increase in correct codes through constrained decoding for full (left) and argument completion (right)
Figure 4 shows key performance metrics for a selection of models, now equipped with constrained decoding. Additionally, Figure 5 shows the rate of correct codes (t) with and without constrained decoding.
Full completion
Constrained decoding increases the correctness across metrics and models. Code Llama (70B) remains the strongest open-source model at 46% correct codes (t). The relative correctness gains through constrained decoding vary strongly between the models but are generally high, at an average of +90% and up to +200% for Code Llama (13B). Qwen2.5-Coder and Llama 3.1, which were previously unable to generate any executable code, now achieve executability rates similar to the other models, while there are slightly more non-executable codes overall. Notably, the number of illegal URLs and HTTP methods is reduced to zero.
Argument completion
In contrast to the unconstrained baseline, where the performance difference across metrics between the full and argument completion setups was not very pronounced, we observe a much more distinct jump in performance when employing constrained decoding as well as a reduced variance between the models. The relative gains in correct codes (t) average at +135% and go up to +879% for Llama 3.1 (8B). Notably, constrained decoding reduces the number of illegal arguments to zero for all models.
@@ 5.1. Results
Table 5. Evaluation results for full completion with constrained decoding
CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Llama 3.1 (8B) Code Llama (13B) Code Llama (70B)
Correct implementations (t) 000.13 000.15 000.40 000.29 000.27 000.07 000.19 000.25 000.46
Correct URLs (e) 000.36 000.56 000.67 000.66 000.56 000.26 000.40 000.60 000.75
Illegal URLs (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00
Correct methods (e) 000.86 000.89 000.93 000.93 000.92 000.40 000.70 000.92 000.93
Illegal methods (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00
Mean argument precision (e) 000.64 000.76 000.78 000.79 000.77 000.58 000.65 000.76 000.86
Mean argument recall (e) 000.66 000.70 000.79 000.77 000.73 000.64 000.66 000.78 000.84
Mean arg. val. conditional accuracy (e) 000.82 000.71 000.94 000.94 000.94 000.74 000.88 000.81 000.96
@@
Table 6. Evaluation results for argument completion with constrained decoding
CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Llama 3.1 (8B) Code Llama (13B) Code Llama (70B)
Correct implementations (t) 000.35 000.31 000.49 000.47 000.48 000.46 000.47 000.44 000.63
Illegal implementations (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00
Mean argument precision (e) 000.90 000.95 000.92 000.92 000.93 000.93 000.91 000.93 000.96
Mean argument recall (e) 000.86 000.90 000.87 000.90 000.89 000.93 000.89 000.94 000.93
Illegal arguments (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00
Mean arg. val. conditional accuracy (e) 000.89 000.81 000.93 000.96 000.96 000.84 000.93 000.86 000.97
5.2. Discussion
In summary, constrained decoding significantly improves the correctness of generated API invocations by eliminating illegal (i.e., hallucinated) URLs, methods, and arguments—a crucial advancement toward reliable generation of API invocation code. The better efficacy for argument completion is expected, as fixing URLs and methods allows the constraints to substantially narrow down the search space. This suggests that constrained decoding may be especially well suited for in-IDE code completion tasks, where parts of the code are already provided by the user.
Models of all sizes benefit from constrained decoding; it especially shapes up to be an expedient way to make mid-size LLMs competitive to larger ones or to make open-source models viable alternatives to commercial closed-source models (e.g., Code Llama (70B) with constrained decoding matches the performance of GPT-4o mini). However, it remains to be investigated why the quantitative benefit provided by constrained decoding is so model-dependent.
In general, we believe that constrained decoding has the potential to be a powerful tool for reducing specific types of errors in code model outputs. It effectively prunes the model’s search space using background knowledge, particularly valuable for complex, requirement-heavy tasks. A key advantage is that constrained decoding guarantees compliance with specified constraints, whereas unconstrained models may inconsistently recall critical information from their training data.
RQ2: How does constrained decoding influence the correctness of LLM-generated web API invocation code? Answer: Constrained decoding is an effective method to improve the overall correctness and reliability of generated API invocations. It provides relative correctness gains of +90% and +135% on average, depending on the setup. Crucially, it guarantees the absence of illegal URLs, HTTP methods, and arguments.
6. Limitations and Threats to Validity
@@
1) Real-world API integration code is diverse and complex. To be able to evaluate API invocations automatically, we had to apply some restrictions and simplifications when creating the tasks in our dataset, as described in Section 2.1. For instance, there is limited program context before the API invocation. Additionally, we evaluate only the correctness of the outgoing request, not how the incoming response is handled. These limitations do, however, not interfere with our investigation of LLM’s ability to recall information about correct API usage—if this information is memorized, they should be able to recall it in our evaluation as well as real-world settings.
@@
2) The quality of the dataset, evaluation, and constraints depends on the quality of the API specifications. We encountered and manually fixed several errors in the real-world specifications. Moreover, the specifications sometimes contain usage constraints that are only explained in a parameter’s free-form textual description and thus not captured by our automated correctness analysis or constraint generator. Future work could use an NLP-based approach to extract these kinds of constraints.
@@
4) Our constraints act only locally from the beginning of an API invocation to its end. When (instead of literals) variables are used as parameter values, we have no control over such a variable’s value or data type. To overcome this, we would need to extend our approach with a backtracking mechanism, which comes with its own set of issues (Jiang et al., 2025).
5) Our implementation requires the user to provide the name and specification of the API, for which invocations are to be generated. The limitation could be addressed by automatically inferring the API referred to in a task description and retrieving the corresponding specification. However, our work targets use cases in the API integrations industry, where usually the customer knows which APIs they want to be integrated, and the developer only has to decide which endpoints to use in order to connect the APIs in the desired way. In this scenario, having to provide the correct API specification to our system is not a real limitation. Our work is not intended for highly exploratory development scenarios, but future work could fill this gap.
6) The executability rate of the code generated by constrained models is slightly lower than that of unconstrained models, which can be attributed to misalignment between the LLMs’ subword tokens and the constraints. This can lead a model to generate tokens it would not have chosen by itself, potentially causing erratic behavior by distorting the model’s next token probability distribution (Beurer-Kellner et al., 2024). This issue can be resolved by switching to a minimally invasive constrained decoding engine (Beurer-Kellner et al., 2024; Ugare et al., 2025). However, sporadic cases of non-executable codes are no big concern for practical applications like IDEs, where code suggestions can be generated in the background and discarded or regenerated if they fail a syntax check.
7. Related Work
@@
Since APIs are such a universal concept, it is important to distinguish between different settings in which APIs are being used and understand their different characteristics when discussing related work. We divide APIs roughly into the following categories: (1) general web APIs(this work), (2) domain-specific web APIs, (3) SDK-wrapped web APIs, (4) local function APIs, and (5) tool APIs for AI agents.
@@
2) Su et al. (Su et al., 2017, 2018) and Hosseini et al. (Hosseini et al., 2021) investigated natural language interfaces to domain-specific web APIs following the Open Data Protocol [^15] https://docs.oasis-open.org/odata/odata/v4.0/odata-v4.0-part1-protocol.html (OData). Their work has a narrow focus on data-centric applications (mostly within the Microsoft ecosystem) and an SQL-like syntax. In contrast, our benchmark targets OpenAPI, a very general and flexible API standard, making the task more open-ended and challenging. Also, instead of ad-hoc API queries by a user, we strive for production-ready API integration code implemented in JavaScript.
3) Works like Dialog2API (Shu et al., 2022) and CloudAPIBench (Jain et al., 2025) focused on SDK-wrapped web API invocations (e.g., for AWS), treating web APIs like local libraries. Both approaches do not generalize due to SDK heterogeneity. In contrast, our benchmark relies on a unified RESTful interface for any web API. This eliminates SDK dependency, enabling direct assessment of models’ core API comprehension rather than SDK-specific knowledge. By standardizing interactions across diverse APIs, we remove evaluation biases and isolate model capabilities from implementation quirks of individual SDKs.
4) Local function APIs are the most common kind of API in typical codebases and studied in many works, e.g., (Zan et al., 2025; Eghbali and Pradel, 2024; Dutta et al., 2024). Invoking a function requires specifying the function name and one argument list. In contrast, general web API invocations are subject to more complex syntax and semantics than such local function calls, as already mentioned in Section 1 (Scope).
5) There is a large body of work concerned with enabling LLM-based AI agents to use tools to interact with their environment (Yehudai et al., 2025). These tools range from simple utilities like calculators (Schick et al., 2023) to web services that enable AI agents to perform tasks such as weather queries or restaurant booking (Song et al., 2023; Li et al., 2023b; Qin et al., 2024). At first sight, this body of work seems related to research on automatic generation of integration code, as tools typically expose a function-like interface. However, both input and output modalities differ between these settings, as well as their non-functional requirements, so benchmarks are not directly transferable to other kinds of API invocations.
@@
Since functional abstractions are essential in programming, virtually every code dataset includes (function) API invocations. Code datasets involving web APIs can be found in the fields of natural language interfaces to web APIs (Su et al., 2017; Hosseini et al., 2021), SDK-based API invocations (Shu et al., 2022; Jain et al., 2025), and tool-using AI agents (Patil et al., 2024; Song et al., 2023; Li et al., 2023b; Qin et al., 2024; Yan et al., 2024). We are not aware of datasets specifically about generating API integration code to invoke REST APIs like we do.
@@
The correctness of AI-generated code is traditionally measured either using exact matches or syntax-based similarity, which both correlate weakly with actual correctness (Evtikhiev et al., 2023), or using functional testing (Chen et al., 2021), which is not directly transferable to web API invocations, as discussed in Section 2.3. This is why we developed a novel approach described in Section 2. Evaluation methods in related works on API invocations mostly rely on variations of the aforementioned standard approaches. Some works use humans (Ding et al., 2025) or LLMs (Qin et al., 2024) as judges, and others execute the API invocation and verify the state of the environment (Shu et al., 2022; Yan et al., 2024). We are not aware of an evaluation method able to provide results as fine-grained as ours.
@@
LLMs display a remarkable capacity to memorize information from their training data—which can be beneficial when it comes to generating factual responses, but also a problem when it comes to leaking private information from the training data. Studies have shown that LLMs memorize, among other things, method definitions and method calls (Yang et al., 2024), but also that they struggle to suggest the correct APIs (Zhuo et al., 2023). These findings are in line with our observation that LLMs seem to know many endpoint URLs and argument names, but are often unable to assemble them to a specification-compliant API invocation that solves the task in question.
If LLMs are unable to recall memorized facts, they tend to hallucinate information that is not grounded in reality. This is a common error of LLMs, and studies have shown that it also exists in code generation, where, among other things, models may hallucinate function names and arguments (Liu et al., 2024b; Dou et al., 2024; Tambon et al., 2025). Our study confirms that these findings also hold for web API endpoints and their arguments, where a high amount of hallucinations can be observed.
@@
Correct usage of APIs is challenging also for humans, as indicated by several studies on API usability issues (Robillard, 2009; Rauf et al., 2019) and real-world API misuses (Bae et al., 2014; Li et al., 2021; Wang and Zhao, 2023). Several (static analysis) methods have been proposed for detecting (and mitigating) API misuses, e.g., (Bae et al., 2014; Binder et al., 2022; Wang and Zhao, 2023). They are, however, primarily designed to aid humans, not LLMs. Using static analyses for evaluation purposes is generally conceivable, but they can only detect whether an API invocation is legal or illegal, not whether it also solves the given task.
Constrained Decoding for Reliable Code Generation
Constrained decoding for code generation commonly enforces syntactic validity using general-purpose constrained decoding frameworks (Willard and Louf, 2023; Beurer-Kellner et al., 2024; Ugare et al., 2025; Dong et al., 2024). These frameworks guide the generation based on regular expressions or context-free grammars, and many of them leverage precomputed finite-state machines to optimize performance. Since syntax errors make up only a small part of errors in LLM-generated code (Dou et al., 2024; Tambon et al., 2025), several works have explored constrained decoding to prevent more advanced kinds of errors. PICARD (Scholak et al., 2021) constrains SQL statements to use only existing tables and columns. Synchromesh (Poesia et al., 2022) uses hand-crafted semantic constraints to make generated DSL code obey certain constrains, e.g., related to scopes and types. Another application (Mündler et al., 2025) uses rules from the TypeScript type system to ensure that generated code is correctly typed. While our constraints do not incorporate a complete type system, we do ensure that arguments passed as literals match the type given in the OpenAPI specification.
Several works have created constraints specifically to mitigate hallucinations related to function/method invocations, which appears to be a common issue (Dou et al., 2024; Tambon et al., 2025). Monitor-Guided Decoding (MGD) (Agrawal et al., 2023) and Repilot (Wei et al., 2023) use constraints provided by a Java language server to prevent various errors, including the hallucination of methods dereferenced from an object. MGD additionally constrains the number of method arguments. MARIN (Chen et al., 2025) mines dependencies in Java programs to combine constrained decoding and RAG at method call sites, preventing hallucinations of methods and their arguments. The method calls targeted by these three works have a much simpler structure than web API invocations targeted by our work, as already discussed at the beginning of this section. Moreover, language servers and dependency mining can only leverage information contained within the code base, but web APIs are specified externally. These approaches are thus in general unsuitable for our use case. MGD shares one feature with our approach in that it also supports multiple constraints (so-called monitors) with individual start and stop conditions. We go, however, one step further and facilitate the interplay between different constraints for smoother transitions between unconstrained and constrained decoding as described in Section 4.
Related works from the line of research on AI agents have applied constrained decoding to ensure that tool APIs are used correctly. ToolDec (Zhang et al., 2023) uses constrained decoding for tools based on Python functions and REST APIs. However, tools are invoked via a simple JSON-like syntax. FANTASE (Wang et al., 2024) combines constrained decoding of tool names and arguments with RAG (Lewis et al., 2020) and a ranked beam search of candidate tool invocations. Invocations use Python syntax. Compared to these works, our work deals with structurally more complex API invocations, necessitating more sophisticated constraint formulations.
Alternative Approaches to Support LLMs during API Invocations
Besides constrained decoding, several other approaches have been proposed to mitigate problems LLMs face when generating API invocations. Most of these works focus on local libraries (primarily in Python), but some also consider web APIs, even though they are not as complex as the ones targeted by our work (cf. discussion at the beginning of this section). These approaches include iterative refinement (Eghbali and Pradel, 2024; Ding et al., 2025), retrieval-augmented generation (RAG) (Patil et al., 2024; Zan et al., 2025; Zhou et al., 2023; Jain et al., 2025), and fine-tuning (Patil et al., 2024; Zan et al., 2025; Dutta et al., 2024) to reduce hallucinations and improve correctness. The most important difference between these previous works and ours is that none of them—iterative refinement, RAG, and fine-tuning—can provide guarantees (additional disadvantages, such as higher computation cost, prompt bloat, and outdated training data, are discussed in Section 1). In contrast, our approach based on constrained decoding guarantees the absence of illegal endpoint URLs, HTTP methods, and request arguments, thereby improving reliability. Pointing out constrained decoding as a powerful part of the solution is an important contribution of our work. However, we believe that all the approaches discussed here should be seen as complementary, promising synergistic effects when combined.
8. Conclusion and Future Work